A borrower takes out a 10-year loan of 1000 at 8% compounded quarterly. The borrower makes quarterly payments. Calculate the amount of principal paid in the ninth payment.
Answer #1

Calculating Quarterly Payment,

using TVM calculation,

PMT = [PV = 1,000, FV = 0, N = 40, I = 0.08/4]

PMT = $36.56

Calculating Loan Balance after 8th Payment,

FV = [PV = 1,000, PMT = -36.56, N = 8, I = 0.08/4]

FV = $857.87

Interest Payment in 9th Payment = (0.08/4)(857.87) = $17.16

Principal Payment in 9th Payment = 36.56 - 17.16

Principal Payment in 9th Payment = $19.40